OpenStudy (roberts.spurs19):

Please help! very urgent! Find the current through and potential difference across each of two 5Ω resistors when they are placed in a circuit containing a 5V battery, and are wired: a) in series, b) in parallel. Please explain how to do each part so I know how to answer similar questions in the future! Thanks

OpenStudy (afifarahman):

\[\frac{ 1 }{R } = \frac{ 1 }{ R _{1} } +\frac{ 1 }{ R _{2} }\]

OpenStudy (afifarahman):

where \[R _{1} = 5 , R _{2} = 5\] Find the value of R and substitute it in equation \[V=IR\]

OpenStudy (afifarahman):

or \[I= \frac{ V }{ R }\] where V=5V

OpenStudy (afifarahman):

same procedure for series. Formula for series is \[R= R _{1} + R _{2}\]

OpenStudy (radar):